• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 19:27:30 on November17,2014

    Entry number 404071

    keyword=Left HRS at 1.1981 GeV/c

    The controls for the left HRS dipole are still not working. The left dipole current is at 397.44 A. Jack was able to get the NMR to lock. This corresponds to a momentum of 1.1981 GeV/c. We are stuck at this momentum until tomorrow morning. At least the
    magnets are matched at the same setting.

    FIGURE 1

    FIGURE 2